# MemoNav: Working Memory Model for Visual Navigation

### v1

@@ 1 Introduction
Nevertheless, existing topological map-based methods still suffer from two major limitations: (a) Unawareness of useful nodes. These methods typically use all node features for generating actions without considering the contribution of each node, thus being easily misled by redundant nodes that are uninformative of the goal. (b) Local representation. Each node feature only represents a small local area in a large scene, limiting the agent’s capacity to learn a higher-level semantic and geometric representation of the entire scene.
To address the aforementioned limitations, we present a novel ImageNav method named MemoNav (refer to Fig. 1), which draws inspiration from the classical concept of working memory in cognitive neuroscience [15] and loosely aligns with the working memory model in human navigation [7].
@@
Consequently, with the synergy of the three representations, MemoNav noticeably outperforms state-of-the-art methods in the Gibson scenes [41], enjoying substantial improvements on multi-goal navigation tasks. Comparison in the Matterport3D scenes [9] further highlights MemoNav’s generalizability.
@@ 2 Related Work
Memory models for reinforcement learning. Several studies [32, 24, 36, 25] draw inspiration from memory mechanisms of the human brain and design reinforcement learning models for reasoning over long time horizons. Ritter et al. [32] proposed an episodic memory storing state transitions for navigation tasks. Lampinen et al. [24] presented hierarchical attention memory as a form of “mental time-travel” [37], which means recovering goal-oriented information from past experiences. Unlike this method, our model retains such information via a novel forgetting module. Expire-span [36] predicts life spans for each memory fragment and permanently deletes expired ones. Our model is different from this work in that we restore forgotten memory if the agent returns to visited places. [25] shares a similar idea but just solves simple 2D grid-world tasks [14] and its memory capacity is fixed. In contrast, our method employs an adaptive working memory to tackle more complicated long-horizon navigation tasks.
@@ 4.1 Selective Forgetting Module
The forgetting module retains a fraction of STM according to the attention scores $\{\alpha_{i}\}_{i=1}^{N_{t}}$ in $\mathcal{D}_{goal}$. These scores reflect the extent to which the goal embedding ${\bm{e}}_{goal}$ attends to each STM feature. After $\mathcal{D}_{goal}$ finishes decoding, the agent temporarily “forgets” a fraction of nodes whose scores rank below a predefined percentage $p$, meaning these nodes are disconnected from their neighbors and excluded from the navigation pipeline in subsequent steps. If the agent returns to a forgotten node, this node will be re-added to the map and processed by the pipeline again. In multi-goal tasks, once a goal is reached, all forgotten nodes will be restored for potential usefulness in locating the next goal. The forgetting module operates in a plug-and-play manner, which means it is not activated during training but switched on during evaluation and deployment (refer to Fig. 7 for visual details). $p$ is set as 20% as we empirically find that this suits most tasks. With this module, the agent can selectively retain the informative STM, while avoiding misleading experiences.
@@ 4.2 Long-Term Memory Generation
The LTM offers two key benefits: it learns a scene-level feature and facilitates feature fusion. A recent study [31] suggests that embodied agents benefit from higher-level 3D environment representations to mitigate partial observability from limited field-of-view sensors. From this viewpoint, the LTM stores a high-level scene representation by aggregating local node features. Moreover, the LTM facilitates feature fusion, especially useful when the topological map is segmented into isolated sub-graphs due to the removal of forgotten nodes. By connecting to every node, the LTM acts as a bypath that aids in feature fusion across these sub-graphs (wherein sub-graphs can broadcast features via 2-hop graph convolution through the LTM node).
@@ 5.2 Compared Methods and Training Details
We follow the training pipeline in [23] to reproduce CNNLSTM and train our MemoNav. These methods are first trained via imitation learning, minimizing the negative log-likelihood of ground-truth actions. Next, the agents are further fine-tuned with proximal policy optimization (PPO) [35] to enhance exploratory ability. The reward setting and auxiliary losses remain the same as in VGM. The evaluation results for ANS and NTS are borrowed from the VGM paper [23] [^3] ANS is designed for exploration while NTS is not open-sourced, so it is not straightforward to reproduce them for multi-goal tasks.. All methods are equipped with a panoramic camera.
@@ 5.6 Visualization of MemoNav
We plot example trajectories of MemoNav in multi-goal tasks, as shown in Fig. 12 of the appendix. These examples show that MemoNav utilizes the adaptive WM to plan efficient paths in challenging multi-goal navigation. For instance, in the 3-goal trajectory, MemoNav strategically forgets the distant topmost node during the first goal and similarly neglects the goal-irrelevant bottom-left nodes in subsequent goals.
To provide a comprehensive overview, we also analyze MemoNav’s limitations by examining its failure episodes in Fig. 13 of the appendix. These failures predominantly fall into four categories: Stopping mistakenly, Missing the goal, Not close enough, and Over-exploring. This categorization not only helps in understanding the current limitations of MemoNav but also guides potential improvements in future iterations.
@@ 6 Conclusion
This paper proposes MemoNav, a novel memory model for ImageNav. This model flexibly retains the informative fraction of the short-term navigation memory via a forgetting module. We also introduce an extra global node as long-term memory to learn a scene-level representation. The retained short-term memory and the long-term memory are encoded by a graph attention module to generate the working memory that is used for generating action. The experimental results show that the MemoNav outperforms previous methods in multi-goal tasks and plans more efficient routes.
@@ 8       Forgotten indices $i_{forgotten}\leftarrow i\left[0\colon n\right]$;
       $\mathbf{\alpha}\leftarrow\mathcal{D}_{goal}.\operatorname{GetAttScores}()$ // Step 4: Action generation

### v2

@@ 1 Introduction
Nevertheless, existing topological map-based methods still suffer from two major limitations: (a) Unawareness of useful nodes. These methods typically use all node features for generating actions without considering the contribution of each node, thus being easily misled by redundant nodes that are uninformative of the goal. (b) Local representation. Each node feature only represents a small area in a large scene, limiting the agent’s capacity to learn a higher-level semantic and geometric representation of the entire scene.
To address these limitations, we present a novel ImageNav method named MemoNav (refer to Fig. 1), which draws inspiration from the classical concept of working memory in cognitive neuroscience [15] and loosely aligns with the working memory model in human navigation [7].
@@
Consequently, with the synergy of the three representations, MemoNav outperforms state-of-the-art methods in the Gibson scenes [42], enjoying substantial improvements on multi-goal navigation tasks. Comparison in the Matterport3D scenes [9] also highlights MemoNav’s superiority.
@@ 2 Related Work
Memory models for reinforcement learning. Several studies [31, 45, 33, 24, 37, 25] draw inspiration from memory mechanisms of the human brain and design reinforcement learning models for reasoning over long time horizons. Ritter et al. [33] proposed an episodic memory storing state transitions for navigation tasks. Lampinen et al. [24] presented hierarchical attention memory as a form of “mental time-travel” [38], which means recovering goal-oriented information from past experiences. Unlike this method, our model retains such information via a novel forgetting module. Expire-span [37] predicts life spans for each memory fragment and permanently deletes expired ones. Our model is different from this work in that we restore forgotten memory if the agent returns to visited places. [25] shares a similar idea but just solves simple 2D grid-world tasks [14] and its memory capacity is fixed. In contrast, our method employs an adaptive working memory to tackle more complicated long-horizon navigation tasks.
@@ 4.1 Selective Forgetting Module
The forgetting module retains a fraction of STM according to the attention scores $\{\alpha_{i}\}_{i=1}^{N_{t}}$ in $\mathcal{D}_{goal}$. These scores reflect the extent to which the goal embedding ${\bm{e}}_{goal}$ attends to each STM feature. After $\mathcal{D}_{goal}$ finishes decoding, the agent temporarily “forgets” a fraction of nodes whose scores rank below a predefined percentage $p$, meaning these nodes are disconnected from their neighbors and excluded from the navigation pipeline in subsequent steps. If the agent returns to a forgotten node, this node will be re-added to the map and processed by the pipeline again. In multi-goal tasks, once a goal is reached, all forgotten nodes will be restored for potential usefulness in locating the next goal. The forgetting module operates in a plug-and-play manner, which means it is not activated during training but switched on during evaluation and deployment (refer to Fig. 7 for details). $p$ is set as 20% as we empirically find that this suits most tasks. With this module, the agent can selectively retain the informative STM, while avoiding misleading experiences.
@@ 4.2 Long-Term Memory Generation
LTM offers two key benefits: it learns a scene-level feature and facilitates feature fusion. A recent study [32] suggests that embodied agents benefit from higher-level environment representations to mitigate partial observability from limited field-of-view sensors. From this viewpoint, the LTM stores a high-level scene representation by aggregating local node features. Moreover, the LTM facilitates feature fusion, especially useful when the topological map is segmented into isolated sub-graphs due to the removal of forgotten nodes. By connecting to every node, the LTM acts as a bypath aiding in feature fusion across these sub-graphs.
@@ 5.2 Compared Methods and Training Details
We follow the training pipeline in [23] to reproduce CNNLSTM and train our MemoNav. These methods are first trained via imitation learning and then further fine-tuned with proximal policy optimization (PPO) [36] (details in the appendix). The evaluation results for ANS and NTS are borrowed from the VGM paper [23] [^3] ANS is designed for exploration while NTS is not open-sourced, so it is not straightforward to reproduce them for multi-goal tasks.. All methods are equipped with a panoramic camera.
@@ 5.6 Visualization of MemoNav
We plot example trajectories of MemoNav in multi-goal tasks, as shown in Fig. 12. These examples show that MemoNav utilizes the adaptive WM to plan efficient paths in challenging multi-goal navigation. For instance, in the 3-goal trajectory, MemoNav strategically forgets the distant topmost node during the first goal and similarly neglects the goal-irrelevant bottom-left nodes in subsequent goals.
To provide a comprehensive overview, we also analyze MemoNav’s limitations by examining its failure episodes in Fig. 13 of the appendix. These failures predominantly fall into four categories: Stopping mistakenly, Missing the goal, Not close enough, and Over-exploring. This analysis not only helps in understanding the limitations of MemoNav but also guides potential improvements in future iterations.
@@ 6 Conclusion
This paper proposes MemoNav, a novel memory model for ImageNav. This model flexibly retains informative short-term navigation memory via a forgetting module. We also introduce an extra global node as long-term memory to learn a scene-level representation. The retained short-term memory and the long-term memory are encoded by a graph attention module to generate the working memory that is used for generating action. The experimental results show that the MemoNav outperforms previous methods in multi-goal tasks and plans more efficient routes.
7 Acknowledgments
This work was supported in part by the National Key R&D Program of China (No. 2022ZD0160102, No. 2022ZD0160100), the National Natural Science Foundation of China (No. U21B2042, No. 62072457), and in part by the 2035 Innovation Program of CAS.
@@ 10      Forgotten indices $i_{forgotten}\leftarrow i\left[0\colon n\right]$;
16       $\mathbf{\alpha}\leftarrow\mathcal{D}_{goal}.\operatorname{GetAttScores}()$
      // Step 4: Action generation
@@ 9.1 Implementation of MemoNav
We follow the training pipeline in [23] to reproduce CNNLSTM and train our MemoNav. These methods are first trained via imitation learning, minimizing the negative log-likelihood of ground-truth actions. Next, the agents are further fine-tuned with PPO [36] to enhance exploratory ability. The reward setting and auxiliary losses remain the same as in VGM.The reward setting and auxiliary losses remain the same as in VGM.